# Numerical Ergodicity of Stochastic Allen–Cahn Equation driven by Multiplicative White Noise

### v1

@@ Abstract.
We establish the unique ergodicity of a fully discrete scheme for monotone SPDEs with polynomial growth drift and bounded diffusion coefficients driven by multiplicative white noise. The main ingredient of our method depends on the satisfaction of a Lyapunov condition followed by a uniform moments’ estimate, combined with the irreducibility and the strong Feller property for full discretization. We transform the original stochastic equation into an equivalent random equation where the discrete stochastic convolutions are uniformly controlled to derive the desired uniform moments’ estimate. Applying the main result to the stochastic Allen–Cahn equation driven by multiplicative white noise indicates that this full discretization is uniquely ergodic for any interface thickness. Numerical experiments validate our theoretical results.
@@
Primary 60H35; 60H15, 65M60
The author is supported by the National Natural Science Foundation of China, No. 12101296, Basic and Applied Basic Research Foundation of Guangdong Province, No. 2024A1515012348, and Shenzhen Basic Research Special Project (Natural Science Foundation) Basic Research (General Project), No. JCYJ20220530112814033.
@@ 1. Introduction
In the settings of the infinite-dimensional case, we note that most of the existing literature focuses on the numerical ergodicity of Lipschitz SPDEs driven by additive white noise or monotone SPDEs driven by trace-class noise; the case of super-linear SPDEs driven by multiplicative white noise is more subtle and challenging. Our main aim is to indicate that the widely studied DIEG scheme (see (die-g)) applied to second-order monotone SPDEs with polynomial growth coefficients driven by nondegenerate multiplicative white noise is uniquely ergodic (see Theorem 3.1). Applying this result to the 1D stochastic Allen–Cahn equation driven by nondegenerate multiplicative white noise indicates that its DIEG scheme is uniquely ergodic for any interface thickness (see Theorem 3.2).
The paper is organized as follows. Section 2 gives the principal assumptions on the considered SPDE. In this part, we show the unique solvability and required properties of the DIEG scheme. The Lyapunov structure, irreducibility, and strong Feller property of the DIEG scheme and application to the stochastic Allen–Cahn equation are explored in Section 3. The theoretical results are verified by numerical experiments in Section 4.
@@ 2.1. Main Assumptions
Denote by $\Delta$ the Dirichlet Laplacian operator on $H$. Then $-\Delta$ possesses a sequence of positve eigenvalues $\{\lambda_{k}\}_{k\in\mathbb{N}_{+}}$ in an increasing order corresponding to the eigenvectors $\{e_{k}\}_{k\in\mathbb{N}_{+}}$ which vanish on the boundary of $(0,1)$, respectively:
@@
(2.9) ${}_{1}\langle x-y,F(x)-F(y)\rangle_{-1}$ $\leq K\|x-y\|^{2},\quad x,y\in\dot{H}^{1},$
@@ 2.2. DIEG scheme and solvability
$\langle A_{N}u^{N},V_{N}\rangle$ $=-\langle\nabla u^{N},\nabla V_{N}\rangle,\quad u^{N},V_{N}\in V_{N},$
@@
(die-g) $X^{N}_{j+1}=X^{N}_{j}+\tau\Delta_{N}X^{N}_{j+1}+\tau\mathcal{P}_{N}F(X^{N}_{j+1})+\mathcal{P}_{N}G(X^{N}_{j})\delta_{j}W,$
@@
It is clear that the DIEG scheme (die-g) is equivalent to the scheme
@@
where $S_{N,\tau}:=({\rm Id}-\tau\Delta_{N})^{-1}$ is a space-time approximation of the continuous semigroup $\{S(t)=e^{\Delta t}:t\geq 0\}$ in one step. Here and what follows, ${\rm Id}$ denotes the identity operator in $V_{N}$. Iterating (2.14) for $m$-times, we obtain
@@
To investigate the solvability of the DIEG scheme (die-g), we need to consider the implicit operator $\hat{F}:V_{N}\to V_{N}$ defined by
@@
Under the condition (2.5) with $(K_{1}-\lambda_{1})\tau<1$, $\hat{F}:V_{N}\to V_{N}$ defined in (2.16) is bijective so that the DIE scheme (die-g) can be uniquely solved pathwise. Morevoer, $\hat{F}$ is an open map, i.e. for each open set $A\in\mathcal{B}(V_{N})$, $\hat{F}(A)$ is also an open set in $\mathcal{B}(V_{N})$.
@@
As $(K_{1}-\lambda_{1})\tau<1$, $\hat{F}$ defined in (2.16) is strictly monotone in the finite-dimensional Hilbert space $V_{N}$ and thus invertible (see, e.g., [19, Theorem C.2]), so that (die-g) is uniquely solved pathwise.
@@ 2.3. Preliminaries on Ergodicity of Markov Chains
Denote by $P:V_{N}\times\mathcal{B}(V_{N})\to[0,1]$ the transition kernel of the Markov chain $\{X_{n}^{N}:n\in\mathbb{N}\}$ generated by (die-g), i.e.,
@@
$\int_{N}P\phi(x)\mu({\rm d}x)=\mu(\phi):=\int_{N}\phi(x)\mu({\rm d}x),\quad\forall~{}\phi\in\mathcal{C}_{b}(H).$
@@
(2.21) $\lim_{m\to\infty}\frac{1}{m}\sum_{n=0}^{m}P_{n}\phi(x)=\mu(\phi)\quad\text{in}~{}L^{2}(H;\mu),\quad\forall~{}\phi\in L^{2}(H;\mu).$
@@ 3. Unique Ergodicity of DIEG
This section will show the unique ergodicity of the DIEG scheme (die-g) and then apply it to the stochastic Allen–Cahn equation (2.4) driven by nondegenerate multiplicative white noise.
@@ 3.1. Lyapunov structure of DIEG
We begin with the Lyapunov structure of the DIEG scheme (die-g). To this end, we need the following uniform estimation for the sum $(W^{N}_{j})_{j\in\mathbb{N}_{+}}$ of discrete stochastic convolutions in (2.15), where we define
@@
Let (2.5) and (2.11) hold with $(K_{1}-\lambda_{1})\tau<1$. Then for any $N\in\mathbb{N}_{+}$, $\tau\in(0,1)$ with $(K_{1}-\lambda_{1})\tau<1$, and $\beta\in[0,1/2)$, there exists a constant $C=C(p,\beta,K_{6})$ such that
@@
Using the discrete Burkholder–Davis–Gundy inequality (see, e.g., [17, Lemma 2.2]), we obtain
@@
Now, we can develop the following Lyapunov structure for the DIEG scheme (die-g).
@@
$\mathbb{E}[\|F(W^{N}_{j})\|_{-1}^{2}]$ $\leq C\mathbb{E}[\|F(W^{N}_{j})\|_{L^{1}}^{2}]\leq C(1+\mathbb{E}[\||W^{N}_{j}(\xi)\|_{L^{q}}^{2q}])$
$\leq C(1+\mathbb{E}[\||W^{N}_{j}(\xi)\|_{\beta}^{2q}])\leq C.$
@@
The result in Theorem 3.1 indicates that $V:V_{N}\to[0,\infty)$ defined by $V(x)=\|x\|^{2}$, $x\in V_{N}$, is a Lyapunov function of the DIEG scheme (die-g).
3.2. Irreducibility and strong Feller property
In this part, we aim to derive the irreducibility and strong Feller property of the transition kernel $P$ defined in (2.20) associated with the DIEG scheme (die-g)
@@ 3.2. Irreducibility and strong Feller property
Let Assumptions 2.1 and 2.2 hold. Then for any $N\in\mathbb{N}_{+}$ and $\tau\in(0,1)$ with $(K_{1}-\lambda_{1})\tau<1$, $\{X_{n}^{N}:n\in\mathbb{N}\}$ is irreducible and strong Feller in $V_{N}$. Consequently, there exists at most, if it exists, one invariant measure of $\{X_{n}^{N}:n\in\mathbb{N}\}$ in $V_{N}$.
@@
as $x+\mathcal{P}_{N}G(x)\delta_{n}W\sim N(x,[\mathcal{P}_{N}G(x)][\mathcal{P}_{N}G(x)]^{\top}\tau)$, where $\mu_{a,b}$ denotes the Gaussian measure in $V_{N}$ with mean $a\in V_{N}$ and variance operator $b\in\mathcal{L}(V_{N})$. It was shown in Lemma 2.1 that $\hat{F}$ is an open map so that $\hat{F}(A)$ is a non-empty open set. Due to the non-degeneracy of $G$ in Assumption 2.2, the Gaussian measure $\mu_{x,[\mathcal{P}_{N}G(x)][\mathcal{P}_{N}G(x)]^{\top}\tau}(\hat{F}(A))$ is non-degenerate. Indeed, for any $v_{N}\in V_{N}\setminus\{0\}$,
@@
as $g$ is invertible in $\mathbb{R}$. It is known that any nondegenerate Gaussian measure in separable Banach space measures any non-empty open set positive, so the open set $\hat{F}(A)\in\mathcal{B}(V_{N})$ has positive measure under $\mu_{x,[\mathcal{P}_{N}G(x)][\mathcal{P}_{N}G(x)]^{\top}\tau}$, which implies $P(x,A)>0$ and shows the irreducibility of $\{X_{n}^{N}:n\in\mathbb{N}\}$ in $V_{N}$.
Now fix $B\in\mathcal{B}(V_{N})$ and let $x,y\in V_{N}$. It follows from (2.18) and Assumption (2.2), respectively, that $\hat{F}$ and $GG^{\top}$ are both continuous. Then, by Lebesgue dominated convergence theorem, we have
$\lim_{y\to x}\mu_{y,[\mathcal{P}_{N}G(y)][\mathcal{P}_{N}G(y)]^{\top}\tau}(\hat{F}(B))=\mu_{x,[\mathcal{P}_{N}G(x)][\mathcal{P}_{N}G(x)]^{\top}\tau}(\hat{F}(B)),$
from which and (3.5) we conclude $\lim_{y\to x}P(y,B)=P(x,B)$. Therefore, $P(\cdot,B)$ is continuous in $V_{N}$ so that $\{X_{n}^{N}:n\in\mathbb{N}\}$ is strong Feller. The uniqueness of the invariant measure, if it exists, of $\{X_{n}^{N}:n\in\mathbb{N}\}$ follows from the irreducibility and strong Feller property. ∎
Now, we can show the unique ergodicity of the DIEG scheme (die-g).
@@
Let Assumptions 2.1 and 2.2 hold with $K_{2}<\lambda_{1}$. Then (die-g) is uniquely ergodic for any $N\in\mathbb{N}_{+}$ and $\tau\in(0,1)$ with $(K_{1}-\lambda_{1})\tau<1$.
@@
The Lyapunov condition (3.3) in Theorem 3.1, together with the Feller property followed from the strong Feller property in Proposition 3.2, imply the existence of an invariant measure for $\{X_{n}^{N}:n\in\mathbb{N}\}$. Combined with the irreducibility and strong Feller property in Proposition 3.2, we conclude the uniqueness of the invariant measure for $\{X_{n}^{N}:n\in\mathbb{N}\}$. ∎
Applying the above result of Theorem 3.1, we have the following unique ergodicity of the DIEG scheme (die-g) applied to the stochastic Allen–Cahn equation (2.4).
@@
Let Assumption 2.2 hold. For any $\epsilon>0$, $N\in\mathbb{N}_{+}$ and $\tau\in(0,1)$ with $(\epsilon^{-2}-\lambda_{1})\tau<1$, (die-g) applied to Eq. (2.4) is uniquely ergodic. ∎
@@
The validity of (2.5) and (2.7) in Assumption 2.1 in the setting of Eq. (2.4) were shown in [13, Section 4]: $K_{1}=\epsilon^{-2}$, $K_{4}=2\epsilon^{-2}$, and $K_{5}=\epsilon^{-2}$. Morevoer, one can take $K_{2}$ to be any negative scalar and thus (2.6) with $K_{2}<\lambda_{1}$ and some $K_{3}>0$; see [15, Theorem 4.4]. ∎

### v2

@@ Abstract.
We establish the unique ergodicity of a fully discrete scheme for monotone SPDEs with polynomial growth drift and bounded diffusion coefficients driven by multiplicative white noise. The main ingredient of our method depends on the satisfaction of a Lyapunov condition followed by a uniform moments’ estimate, combined with the regularity property for the full discretization. We transform the original stochastic equation into an equivalent random equation where the discrete stochastic convolutions are uniformly controlled to derive the desired uniform moments’ estimate. Applying the main result to the stochastic Allen–Cahn equation driven by multiplicative white noise indicates that this full discretization is uniquely ergodic for any interface thickness. Numerical experiments validate our theoretical results.
@@
Primary 60H35; Secondary 60H15, 65M60
The author is supported by the National Natural Science Foundation of China, No. 12101296, Basic and Applied Basic Research Foundation of Guangdong Province, No. 2024A1515012348, and Shenzhen Basic Research Special Project (Natural Science Foundation) Basic Research (General Project), No. JCYJ20220530112814033 and JCYJ20240813094919026.
@@ 1. Introduction
In the settings of the infinite-dimensional case, we note that most of the existing literature focuses on the numerical ergodicity of Lipschitz SPDEs driven by additive white noise or monotone SPDEs driven by trace-class noise; the case of super-linear SPDEs driven by multiplicative white noise is more subtle and challenging. Our main aim is to indicate that the widely studied DIEG scheme (see (DIEG)) applied to second-order monotone SPDEs with polynomial growth coefficients driven by nondegenerate multiplicative white noise is uniquely ergodic (see Theorem 3.1). Applying this result to the 1D stochastic Allen–Cahn equation driven by nondegenerate multiplicative white noise indicates that its DIEG scheme is uniquely ergodic for any interface thickness (see Theorem 3.2).
The paper is organized as follows. Section 2 gives the principal assumptions on the considered SPDE. In this part, we show the unique solvability and required properties of the DIEG scheme. The Lyapunov structure and regularity property of the DIEG scheme with application to the stochastic Allen–Cahn equation are explored in Section 3. The theoretical results are verified by numerical experiments in Section 4.
@@ 2.1. Main Assumptions
Denote by $\Delta$ the Dirichlet Laplacian operator on $H$. Then $-\Delta$ possesses a sequence of blue positive eigenvalues $\{\lambda_{k}\}_{k\in\mathbb{N}_{+}}$ in an increasing order corresponding to the eigenvectors $\{e_{k}\}_{k\in\mathbb{N}_{+}}$ which vanish on the boundary of $(0,1)$, respectively:
@@
Remark 2.1.
It is clear that the corresponding function $f(\xi):=\epsilon^{-2}(\xi-\xi^{3})$, $\xi\in\mathbb{R}$, in the stochastic Allen–Cahn equation (2.4) satisfies Assumption 2.1; see [17, Example 2.1].
@@
(2.9) ${}_{1}\langle x-y,F(x)-F(y)\rangle_{-1}$ $\leq K_{1}\|x-y\|^{2},\quad x,y\in\dot{H}^{1},$
@@ 2.2. DIEG scheme and solvability
$\langle\Delta_{N}u^{N},v_{N}\rangle$ $=-\langle\nabla u^{N},\nabla v_{N}\rangle,\quad u^{N},v_{N}\in V_{N},$
@@
(DIEG) $X^{N}_{j+1}=X^{N}_{j}+\tau\Delta_{N}X^{N}_{j+1}+\tau\mathcal{P}_{N}F(X^{N}_{j+1})+\mathcal{P}_{N}G(X^{N}_{j})\delta_{j}W,$
@@
It is clear that the DIEG scheme (DIEG) is equivalent to the scheme
@@
where $S_{N,\tau}:=({\rm Id}-\tau\Delta_{N})^{-1}$ is a space-time approximation of the continuous semigroup $\{S(t)=e^{\Delta t}:t\geq 0\}$ in one step. Here and in what follows, ${\rm Id}$ denotes the identity operator in $V_{N}$. Iterating (2.14) for $m$-times, we obtain
@@
To investigate the solvability of the DIEG scheme (DIEG), we need to consider the implicit operator $\hat{F}:V_{N}\to V_{N}$ defined by
@@
Under the condition (2.5) with $(K_{1}-\lambda_{1})\tau<1$, $\hat{F}:V_{N}\to V_{N}$ defined in (2.16) is bijective so that the DIEG scheme (DIEG) can be uniquely solved pathwise. Morevoer, $\hat{F}$ is an open map, i.e. for each open set $A\in\mathcal{B}(V_{N})$, $\hat{F}(A)$ is also an open set in $\mathcal{B}(V_{N})$.
@@
As $(K_{1}-\lambda_{1})\tau<1$, $\hat{F}$ defined in (2.16) is strictly monotone in the finite-dimensional Hilbert space $V_{N}$ and thus invertible (see, e.g., [19, Theorem C.2]), so that (DIEG) is uniquely solved pathwise.
@@ 2.3. Preliminaries on Ergodicity of Markov Chains
Denote by $P:V_{N}\times\mathcal{B}(V_{N})\to[0,1]$ the transition kernel of the Markov chain $\{X_{n}^{N}:n\in\mathbb{N}\}$ generated by (DIEG), i.e.,
@@
We also use $P_{n}$, $n\in\mathbb{N}_{+}$, to denote the corresponding Markov semigroup on $\mathcal{B}(V_{N})$:
$P_{n}\phi(x):=\mathbb{E}[\phi(X_{n}^{N})|X_{0}^{N}=x],\quad x\in V_{N},~{}\phi\in\mathcal{B}_{b}(V_{N}).$
@@
$\int_{V_{N}}P\phi(x)\mu({\rm d}x)=\mu(\phi):=\int_{V_{N}}\phi(x)\mu({\rm d}x),\quad\forall~{}\phi\in\mathcal{C}_{b}(H).$
@@
(2.21) $\lim_{m\to\infty}\frac{1}{m}\sum_{n=0}^{m}P_{n}\phi(x)=\mu(\phi)\quad\text{in}~{}L^{2}(V_{N};\mu),\quad\forall~{}\phi\in L^{2}(V_{N};\mu).$
@@ 3. Unique Ergodicity of DIEG
This section will show the unique ergodicity of the DIEG scheme (DIEG) and then apply it to the stochastic Allen–Cahn equation (2.4) driven by nondegenerate multiplicative white noise.
@@ 3.1. Lyapunov structure of DIEG
We begin with the Lyapunov structure of the DIEG scheme (DIEG). To this end, we need the following uniform estimation for the sum $(W^{N}_{j})_{j\in\mathbb{N}_{+}}$ of discrete stochastic convolutions in (2.15), where we define
@@
Let (2.5) and (2.11) hold with $(K_{1}-\lambda_{1})\tau<1$. Then for any $p\geq 2$, $N\in\mathbb{N}_{+}$, $\tau\in(0,1)$ with $(K_{1}-\lambda_{1})\tau<1$, and $\beta\in[0,1/2)$, there exists a positive constant $C=C(p,\beta,K_{6})$ such that
@@
Using the discrete Burkholder–Davis–Gundy (BDG) inequality (see, e.g., [17, Lemma 2.2]), we obtain
@@
Now, we can develop the following Lyapunov structure for the DIEG scheme (DIEG).
@@
$\mathbb{E}[\|F(W^{N}_{j})\|_{-1}^{2}]$ $\leq C\mathbb{E}[\|F(W^{N}_{j})\|_{L^{1}}^{2}]\leq C(1+\mathbb{E}[\||W^{N}_{j}\|_{L^{q}}^{2q}])$
$\leq C(1+\mathbb{E}[\||W^{N}_{j}\|_{\beta}^{2q}])\leq C.$
@@
The result in Proposition 3.1 indicates that $V:V_{N}\to[0,\infty)$ defined by $V(x)=\|x\|^{2}$, $x\in V_{N}$, is a Lyapunov function of the DIEG scheme (DIEG).
Remark 3.2.
We impose the bounded assumption on the diffusion coefficient in the present white noise case. For an unbounded diffusion coefficient, such stochastic-random transform argument would fail as one could not derive the estimate (3.2) for $W^{N}_{j}$ defined in (3.1).
3.2. Regularity Property
In this part, we aim to derive the regularity property of the transition kernel $P$ defined in (2.20) associated with the DIEG scheme (DIEG) in the sense that all transition probabilities of (DIEG) are equivalent.
@@ 3.2. Regularity Property
Let Assumptions 2.1 and 2.2 hold. Then for any $N\in\mathbb{N}_{+}$ and $\tau\in(0,1)$ with $(K_{1}-\lambda_{1})\tau<1$, $P$ is regular. Consequently, there exists at most, if it exists, one invariant measure of $\{X_{n}^{N}:n\in\mathbb{N}\}$ in $V_{N}$.
@@
as $x+\mathcal{P}_{N}G(x)\delta_{n}W\sim N(x,[\mathcal{P}_{N}G(x)][\mathcal{P}_{N}G(x)]^{\top}\tau)$, where $\mu_{a,b}$ denotes the Gaussian measure in $V_{N}$ with mean $a\in V_{N}$ and variance operator $b\in\mathcal{L}(V_{N})$. It was shown in Lemma 2.1 that $\hat{F}$ is an open map so that $\hat{F}(A)$ is a non-empty open set. Due to the non-degeneracy of $G$ in Assumption 2.2, the Gaussian measure $\mu_{x,[\mathcal{P}_{N}G(x)][\mathcal{P}_{N}G(x)]^{\top}\tau}(\hat{F}(A))$ is non-degenerate in $\mathcal{B}(V_{N})$. Indeed, for any $v_{N}\in V_{N}\setminus\{0\}$,
@@
as $g$ is invertible in $\mathbb{R}$. It is known that any nondegenerate Gaussian measure in separable Banach space measures any non-empty open set positive, so the open set $\hat{F}(A)\in\mathcal{B}(V_{N})$ has positive measure under $\mu_{x,[\mathcal{P}_{N}G(x)][\mathcal{P}_{N}G(x)]^{\top}\tau}$, which implies $P(x,A)>0$ and shows the irreducibility of $\{X_{n}^{N}:n\in\mathbb{N}\}$ in $V_{N}$. It is well-known that all non-degenerate Gaussian measures are equivalent in the finite-dimensional case. Thus, $P$ is regular, and by Doob theorem, it possesses at most one invariant measure. ∎
Now, we can show the unique ergodicity of the DIEG scheme (DIEG).
@@
Let Assumptions 2.1 and 2.2 hold with $K_{2}<\lambda_{1}$. Then (DIEG) is uniquely ergodic for any $N\in\mathbb{N}_{+}$ and $\tau\in(0,1)$ with $(K_{1}-\lambda_{1})\tau<1$.
@@
The Lyapunov condition (3.3) in Theorem 3.1, together with the Feller property followed from the regularity property in Proposition 3.2, imply the existence of an invariant measure for $\{X_{n}^{N}:n\in\mathbb{N}\}$. Combined with the regularity property in Proposition 3.2, we conclude the uniqueness of the invariant measure for $\{X_{n}^{N}:n\in\mathbb{N}\}$. ∎
Applying the above result of Theorem 3.1, we have the following unique ergodicity of the DIEG scheme (DIEG) applied to the stochastic Allen–Cahn equation (2.4).
@@
Let Assumption 2.2 hold. For any $\epsilon>0$, $N\in\mathbb{N}_{+}$ and $\tau\in(0,1)$ with $(\epsilon^{-2}-\lambda_{1})\tau<1$, (DIEG) applied to Eq. (2.4) is uniquely ergodic. ∎
@@
The validity of (2.5) and (2.7) in Assumption 2.1 in the setting of Eq. (2.4) were shown in [14, Section 4]: $K_{1}=\epsilon^{-2}$, $K_{4}=2\epsilon^{-2}$, and $K_{5}=\epsilon^{-2}$. Morevoer, one can take $K_{2}$ to be any negative scalar and thus (2.6) with $K_{2}<\lambda_{1}$ and some $K_{3}>0$; see [15, Theorem 4.4]. ∎